# Distance-based measures and Epsilon-measures for measurement-based quantum resources

### v3

@@ II.1 Quantum Measurements and quantum channels
A quantum measurement $M=\{M(x)\in\mathcal{L}(\mathcal{H})\}_{x\in\Omega_{M}}$ acting on Hilbert space $\mathcal{H}$ is defined as a set of positive semidefinite matrices acting on Hilbert space $\mathcal{H}$ such that $\sum_{x\in\Omega_{M}}M(x)=\mathbbm{1}_{\mathcal{H}}$ where $\mathcal{L}(\mathcal{H})$ is the set of all linear operators on Hilbert space $\mathcal{H}$, $\Omega_{M}$ is the outcome set of $M$ and $\mathbbm{1}_{\mathcal{H}}$ is the idenitity matrix on the Hilbert space $\mathcal{H}$ [24]. In this work, we restrict ourselves to finite dimensional Hilbert spaces and measurements with finite number of outcomes. Here we call $M(x)$ as the POVM elements of $M$ for all $x$. The measurement $M$ is said to be projective if $M^{2}(x)=M(x)\forall x\in\Omega_{M}$. When the measurement $M$ is performed on a system with quantum state $\rho\in\mathcal{S}(\mathcal{H})$, the probability of an arbitrary outcome $x$ is $Tr[\rho M(x)]$ where $\mathcal{S}(\mathcal{H})$ is the set of all density matrices on Hilbert space $\mathcal{H}$. We denote the set of all measurements acting on Hilbert space $\mathcal{H}$ as $\mathscr{M}(\mathcal{H})$. Given a pair of quantum measurements $M\in\mathscr{M}(\mathcal{H})$ and $N\in\mathscr{M}(\mathcal{K})$, the measurement $M\otimes N:=\{(M\otimes N)(x,y)=M(x)\otimes N(y)\}_{x\in\Omega_{M},y\in\Omega_{N}}\in\mathscr{M}(\mathcal{H}\otimes\mathcal{K})$. Clearly, the outcome set of the measurement $M\otimes N$ is $\Omega_{M\otimes N}=\Omega_{M}\times\Omega_{N}$. We denote the one-outcome trivial measurement acting on Hilbert space $\mathcal{H}$ as $\mathcal{T}_{\mathcal{H}}:=\{\mathbbm{1}_{\mathcal{H}}\}$. Implementation of $\mathcal{T}_{\mathcal{H}}$ is equivalent to performing “no measurement”. Clearly, $\mathcal{T}_{\mathcal{H}_{1}\otimes\mathcal{H}_{2}}=\mathcal{T}_{\mathcal{H}_{1}}\otimes\mathcal{T}_{\mathcal{H}_{2}}$ If a quantum measurement $M\in\mathscr{M}(\mathcal{H})$ is the probabilistic mixing of a pair of quantum measurements $M_{1}\in\mathscr{M}(\mathcal{H})$ and $M_{2}\in\mathscr{M}(\mathcal{H})$ then $M=\{M(x)=pM_{1}(x)+(1-p)M_{2}(x)\}_{x\in\Omega_{M}}$ where $\Omega_{M}=\Omega_{M_{1}}=\Omega_{M_{2}}$ and $p$ being the probability. We denote it as $M=pM_{1}+(1-p)M_{2}$. We denote a process where a given measurement $M$ (as an input) is probabilistically mixed with a fixed channel $N$ as
@@
$Tr[\Lambda(W)Z]=Tr[W\Lambda^{\dagger}(Z)]\leavevmode\nobreak\ \forall W\in\mathcal{L}(\mathcal{H}),Z\in\mathcal{L}(\mathcal{K}).$ (2)
@@
Superchannels are the physically realizable transformations between quantum channels. A super-channel $\hat{\Xi}$ transforming a quantum channel in $\Lambda\in\mathscr{C}(\mathcal{H}_{1},\mathcal{H}_{2})$ into a quantum channel in $\hat{\Xi}[\Lambda]\in\mathscr{C}(\mathcal{K}_{1},\mathcal{K}_{2})$ can be written as
@@
where the quantum channel $\Theta_{pre}:\mathcal{L}(\mathcal{K}_{1})\rightarrow\mathcal{L}(\mathcal{H}\otimes\mathcal{R})$ is known as pre-processing and the quantum channel , $\Theta_{post}:\mathcal{L}(\mathcal{H}_{2}\otimes\mathcal{R})\rightarrow\mathcal{L}(\mathcal{K}_{2})$ is known as post-processing and $\mathbbm{I}_{\mathcal{R}}$ is the identity quantum channel on $\mathcal{L}(\mathcal{R})$ [25].
@@
where we denote the trace norm by .$\mid\mid.\mid\mid_{1}$, and $\dim(\mathcal{H}_{A})=\dim(\mathcal{H}_{B})$. This distance is known as the diamond distance.
@@
Note that the trace norm of an arbitrary linear map $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})$ is defined as $\mid\mid\Phi\mid\mid_{1}:=\rm{max}_{X\in\mathcal{L}(\mathcal{H})}\{\mid\mid\Phi(X)\mid\mid_{1}\mid\leavevmode\nobreak\ \mid\mid X\mid\mid_{1}\leq 1\}$. Now, the diamond norm of an arbitrary linear map $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})$ is originally defined as $\mid\mid\Phi\mid\mid_{\Diamond}:=\mid\mid\Phi\otimes\mathbbm{I}_{\mathcal{H}_{B}}\mid\mid_{1}$ with $\dim(\mathcal{H})=\dim(\mathcal{H}_{B})$ [26]. But note that $(\Gamma_{M_{1}}-\Gamma_{M_{2}})$ is a Hermiticity-preserving map and we know that for the Hermiticity-preserving map, the original definition coincides with Eq. (5) [26]. Furthermore, one can show that $\mid\mid\Phi\mid\mid_{\Diamond}=\mid\mid\Phi\otimes\mathbbm{I}_{\mathcal{H}_{R}}\mid\mid_{1}$ for any $\mathcal{H}_{R}$ with $\dim(\mathcal{H})\leq\dim(\mathcal{H}_{R})$ [26]. It should be mentioned here that for a pair of arbitrary linear maps $\Phi_{1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{K}_{1}$ and $\Phi_{2}:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{K}_{2}$, the equality $\mid\mid\Phi_{1}\otimes\Phi_{2}\mid\mid_{\Diamond}=\mid\mid\Phi_{1}\mid\mid_{\Diamond}.\mid\mid\Phi_{2}\mid\mid_{\Diamond}$ holds [26]. Furthermore, if $\Phi:\mathcal{L}(\mathcal{H})\otimes\mathcal{L}(\mathcal{K})$ is a quantum channel, $\mid\mid\Phi\mid\mid_{\Diamond}=1$ [26].
@@
After performing a measurement $M=\{M(x)\}$, one can post-process the outcomes which is equivalent to performing another measurement $N=N(y)$. For this purpose, if one uses the probability distribution $\nu_{x}=\{\nu_{x}(y)\}$ to post-process the outcome $x$ then
$N(y)=\sum_{x\in\Omega_{M}}\nu_{x}(y)M(x)\leavevmode\nobreak\ \forall y\in\Omega_{N}.$ (12)
A measurement $M$ can be performed using many different quantum instruments. A quantum instrument $\mathcal{I}$ is defined as a set of CP trace non-increasing linear maps $\{\Phi_{x}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})\}_{x\in\hat{\Omega}_{\mathcal{I}}}$ that sums up to a quantum channel $\Phi$ where $\hat{\Omega}_{\mathcal{I}}$ is the outcome set of the instrument $\mathcal{I}$ [24]. It is well-known that both quantum measurements and quantum channels are special cases of quantum instruments. If the instrument $\mathcal{I}$ implements the measurement $M\in\mathscr{M}(\mathcal{H})$, then
@@
hold for all $x\in\hat{\Omega}_{\mathcal{I}}$ and $\hat{\Omega}_{\mathcal{I}}=\Omega_{M}$.
@@
Clearly, the measurement $\mathcal{I}^{\dagger}[N]$ has the outcome set $\Omega_{\mathcal{I}^{\dagger}[N]}=(\hat{\Omega}_{\mathcal{I}}\times\Omega_{N})$. A general transformation of a measurement can be the compositions of postprocessing of outcomes, probabilistic mixinging with another measurement and action of quantum instruments. Similar to the sets of quantum channels, transformations of a set of measurements is slightly more general. Such transformations allow the controlled implemention of measurements in that set, in general.
@@
$\Gamma_{M}(\rho)=\sum_{a}Tr[\rho M(a)]\ket{a}\bra{a}$ | (15)
for all $\rho\in\mathcal{L}(\mathcal{H}_{A})$ where {|$\{\ket{a}\}$ is a chosen orthonormal basis of Hilbert space $\mathcal{H}_{\Omega_{M}}$ with dimension $|\Omega_{M}|$ where we denote the cardinality of a set by the symbol |.||.|| . |. Clearly, $\Gamma_{M}\in\mathscr{C}(\mathcal{H}_{A},\mathcal{H}_{\Omega_{M}})$ and $\Gamma_{M\otimes N}=\Gamma_{M}\otimes\Gamma_{N}$ for two arbitrary measurements $M\in\mathscr{M}(\mathcal{H}_{A})$ and $N\in\mathscr{M}(\mathcal{H}_{B})$.
@@
$=$ $\sum_{ab}Tr[\rho_{AB}(M(a)\otimes N(b))]\ket{a}\bra{a}$ |
$=$ $\sum_{a}Tr[\rho_{AB}(M(a)\otimes\sum_{b}N(b))]\ket{a}\bra{a}$ |
$=$ $\sum_{a}Tr[\rho_{AB}(M(a)\otimes\mathbbm{1}_{\mathcal{H}_{B}})]\ket{a}\bra{a}$ |
@@
Therefore, $Tr_{\mathcal{H}_{\Omega_{N}}}$ is equivalent to the marginalization operation on the measurement $N$. Clearly, for an arbitrary measurement $A=\{A(a_{1},a_{2})\}\in\mathscr{M}(\mathcal{H})$ with outcome set $\Omega_{A}=X\times Y$ where the set $X$ and $Y$ are some sets of natural numbers and $a_{1}\in X$ and $a_{2}\in Y$ we can define a channel $\Gamma_{A}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$ with $\dim{\mathcal{H}_{X}}=|X|$ and $\dim{\mathcal{H}_{Y}}=|Y|$. Moreover, we denote the merginalization process as $\mathbf{Merg}_{Y}$. In other words, $\mathbf{Merg}_{Y}(A)=\{[\mathbf{Merg}_{Y}(A)](a_{1})=\sum_{a_{2}}A(a_{1},a_{2})\}$. Clearly,
$\Gamma_{\mathbf{Merg}_{Y}(A)}=(\mathbbm{I}_{\mathcal{H}_{X}}\otimes Tr_{\mathcal{H}_{Y}})\circ\Gamma_{A}.$ (17)
@@
Consider two arbitrary measurements $M=\{M(a_{1},a_{2})\}\in\mathscr{M}(\mathcal{H})$ and $N=\{N(b_{1},b_{2})\}\in\mathscr{M}(\mathcal{H})$ with $\Omega_{A}=\Omega_{B}=X\times Y$. Clearly,
@@
where $M_{CBA}=\mathbbm{I}^{\dagger}_{\mathcal{H}_{C}}\otimes\mathbf{SWAP}^{\dagger}_{\mathcal{H}_{B}\leftrightarrow\mathcal{H}_{A}}(M_{CAB})$ and $N_{CBA}=\mathbbm{I}^{\dagger}_{\mathcal{H}_{C}}\otimes\mathbf{SWAP}^{\dagger}_{\mathcal{H}_{B}\leftrightarrow\mathcal{H}_{A}}(N_{CAB})$ In Section III.1, we will define the distance for two sets of measurements.
@@ II.2 Aspects of a resource measure based on set of quantum measurements
In case of measurements-based resource theory, one can make the following assumptions-
@@
    If two objects $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are free if and only if $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is free.
@@
    If two transformations $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ are free if and only if $\mathcal{W}_{1}\otimes\mathcal{W}_{2}$ is free.
@@
The assumption A6 will not be not used everywhere in the manuscript. We will explicitly mention wherever we use the assumption [A6]. Merginalization is a free transformation for the resource theory of measurement incompatibility and the resource theory of measurement coherence. Now, it is known that the post-processing of outcomes is not free operations for the resource theory of measurement sharpness, in general [13, 14]. However, merginalization is very special case of post-processing of outcomes that transforms an arbitrary measurement $M\in\mathscr{M}(\mathcal{H})$ to $\mathcal{T}_{\mathcal{H}}$ which is a trivial object. Therefore, technically, the merginalization should be the free transformation for the the resource theory of measurement sharpness. But in Ref. [14], only the outcome set preserving free transformations for the resource theory of measurement sharpness have been considered.
@@
    (Non-negativity and faithfulness): $\mathbbm{R}(\mathcal{M})\geq 0$ for all objects $\mathcal{M}$ and $\mathbbm{R}(\mathcal{M})=0$ if and only if $\mathcal{M}\in\mathcal{F}$
@@
    (Monotonicity): $\mathbbm{R}(\mathcal{M})\geq\mathbbm{R}(\mathcal{W}(\mathcal{M}))$ for all objects $\mathcal{M}$ and for all free transformations $\mathcal{W}$ i.e., for all $\mathcal{W}\in\mathcal{O}$
@@
    (Convexity): $\mathbbm{R}(\sum^{n}_{i}p_{i}\mathcal{M}_{i})\leq\sum^{n}_{i}p_{i}\mathbbm{R}(\mathcal{M}_{i})$ for all sets of objects $\{\mathcal{M}_{i}\}^{n}_{i=1}$ and all probability distributions $\{p_{i}\}^{n}_{i=1}$
@@
For convex measurement-based resource theories, the resource robustness of an arbitrary set of measurements =$\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
$\begin{split}\mathscr{R}(\mathcal{M})=\min\leavevmode\nobreak\ &\left.r\right.\\ \text{s.t.}\leavevmode\nobreak\ &\left.\mathcal{N}=\{N_{i}=\frac{M_{i}}{1+r}+\frac{r\tilde{M}_{i}}{1+r}\}\in\mathcal{F}_{\mathcal{H}_{A}}\right.\\ &\left.\tilde{\mathcal{M}}=\{\tilde{M}_{i}\}\subset\mathscr{M}(\mathcal{H}_{A}).\right.\end{split}$ (23)
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the set of measurements $\tilde{\mathcal{M}}$.
Similarly, for convex measurement-based resource theories, the resource weight of an arbitrary set of measurements =$\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
$\begin{split}\mathscr{W}(\mathcal{M})=\min\leavevmode\nobreak\ &\left.r\right.\\ \text{s.t.}\leavevmode\nobreak\ &\left.\mathcal{M}=\{M_{i}=\frac{\tilde{N}_{i}}{1+r}+\frac{r\tilde{M}_{i}}{1+r}\}\right.\\ &\left.\tilde{\mathcal{M}}=\{\tilde{M}_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})\right.\\ &\left.\tilde{\mathcal{N}}=\{\tilde{N}_{i}\}\in\mathcal{F}_{\mathcal{H}_{A}}.\right.\end{split}$ (24)
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the set of measurements $\tilde{\mathcal{M}}$.
@@
For any resource measure $\mathbbm{R}$ in any measurement-based resource theory, and for a generic distance $\mathbf{D}$, the corresponding $\epsilon$-inf-measure $\mathbbm{R}^{\mathbf{D}}_{inf,\varepsilon}$ of a set of measurements =$\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H})$ is defined as
$\mathbbm{R}^{\mathbf{D}}_{inf,\epsilon}(\mathcal{M})=\inf_{\begin{subarray}{c}\mathcal{N}\subset\mathscr{M}(\mathcal{H}),\\ \mathbf{D}(\mathcal{M},\mathcal{N})\leq\epsilon,\leavevmode\nobreak\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N}).$ (25)

### v5

@@ II.1 Quantum Measurements and quantum channels
A quantum measurement $M=\{M(x)\in\mathcal{L}(\mathcal{H})\}_{x\in\Omega_{M}}$ acting on Hilbert space $\mathcal{H}$ is defined as a set of positive semidefinite matrices acting on Hilbert space $\mathcal{H}$ such that $\sum_{x\in\Omega_{M}}M(x)=\mathbbm{1}_{\mathcal{H}}$ where $\mathcal{L}(\mathcal{H})$ is the set of all linear operators on Hilbert space $\mathcal{H}$, $\Omega_{M}$ is the outcome set of $M$ and $\mathbbm{1}_{\mathcal{H}}$ is the idenitity matrix on the Hilbert space $\mathcal{H}$ [24]. In this work, we restrict ourselves to finite dimensional Hilbert spaces and measurements with finite number of outcomes. Here we call $M(x)$ as the POVM elements of $M$ for all $x$. The measurement $M$ is said to be projective if $M^{2}(x)=M(x)\forall x\in\Omega_{M}$. When the measurement $M$ is performed on a system with quantum state $\rho\in\mathcal{S}(\mathcal{H})$, the probability of an arbitrary outcome $x$ is $Tr[\rho M(x)]$ where $\mathcal{S}(\mathcal{H})$ is the set of all density matrices on Hilbert space $\mathcal{H}$. We denote the set of all measurements acting on Hilbert space $\mathcal{H}$ as $\mathscr{M}(\mathcal{H})$. Given a pair of quantum measurements $M\in\mathscr{M}(\mathcal{H})$ and $N\in\mathscr{M}(\mathcal{K})$, the measurement $M\otimes N:=\{(M\otimes N)(x,y)=M(x)\otimes N(y)\}_{x\in\Omega_{M},y\in\Omega_{N}}\in\mathscr{M}(\mathcal{H}\otimes\mathcal{K})$. Clearly, the outcome set of the measurement $M\otimes N$ is $\Omega_{M\otimes N}=\Omega_{M}\times\Omega_{N}$. We denote the one-outcome trivial measurement acting on Hilbert space $\mathcal{H}$ as $\mathcal{T}_{\mathcal{H}}:=\{\mathbbm{1}_{\mathcal{H}}\}$. Implementation of $\mathcal{T}_{\mathcal{H}}$ is equivalent to performing ``no measurement". Clearly, $\mathcal{T}_{\mathcal{H}_{1}\otimes\mathcal{H}_{2}}=\mathcal{T}_{\mathcal{H}_{1}}\otimes\mathcal{T}_{\mathcal{H}_{2}}$ If a quantum measurement $M\in\mathscr{M}(\mathcal{H})$ is the probabilistic mixing of a pair of quantum measurements $M_{1}\in\mathscr{M}(\mathcal{H})$ and $M_{2}\in\mathscr{M}(\mathcal{H})$ then $M=\{M(x)=pM_{1}(x)+(1-p)M_{2}(x)\}_{x\in\Omega_{M}}$ where $\Omega_{M}=\Omega_{M_{1}}=\Omega_{M_{2}}$ and $p$ being the probability. We denote it as $M=pM_{1}+(1-p)M_{2}$. We denote a process where a given measurement $M$ (as an input) is probabilistically mixed with a fixed channel $N$ as
@@
$Tr[\Lambda(W)Z]=Tr[W\Lambda^{\dagger}(Z)]\penalty 10000\ \forall W\in\mathcal{L}(\mathcal{H}),Z\in\mathcal{L}(\mathcal{K}).$ (2)
@@
Superchannels are the physically realizable transformations between quantum channels. A super-channel $\hat{\Xi}$ transforming a quantum channel in $\Lambda\in\mathscr{C}(\mathcal{H}_{1},\mathcal{H}_{2})$ into a quantum channel $\hat{\Xi}[\Lambda]\in\mathscr{C}(\mathcal{K}_{1},\mathcal{K}_{2})$ can be written as
@@
where the quantum channel $\Theta_{pre}:\mathcal{L}(\mathcal{K}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{1}\otimes\mathcal{R})$ is known as pre-processing and the quantum channel , $\Theta_{post}:\mathcal{L}(\mathcal{H}_{2}\otimes\mathcal{R})\rightarrow\mathcal{L}(\mathcal{K}_{2})$ is known as post-processing and $\mathbbm{I}_{\mathcal{R}}$ is the identity quantum channel on $\mathcal{L}(\mathcal{R})$ [25].
@@
where we denote the trace norm by $\mid\mid.\mid\mid_{1}$, and $\dim(\mathcal{H}_{A})=\dim(\mathcal{H}_{B})$. This distance is known as the diamond distance.
@@
Note that the trace norm of an arbitrary linear map $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})$ is defined as $\mid\mid\Phi\mid\mid_{1}:=\rm{max}_{X\in\mathcal{L}(\mathcal{H})}\{\mid\mid\Phi(X)\mid\mid_{1}\mid\penalty 10000\ \mid\mid X\mid\mid_{1}\leq 1\}$. Now, the diamond norm of an arbitrary linear map $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})$ is originally defined as $\mid\mid\Phi\mid\mid_{\Diamond}:=\mid\mid\Phi\otimes\mathbbm{I}_{\mathcal{H}_{B}}\mid\mid_{1}$ with $\dim(\mathcal{H})=\dim(\mathcal{H}_{B})$ [26]. But note that $(\Gamma_{M_{1}}-\Gamma_{M_{2}})$ is a Hermiticity-preserving map and we know that for the Hermiticity-preserving map, the original definition coincides with Eq. (5) [26]. Furthermore, one can show that $\mid\mid\Phi\mid\mid_{\Diamond}=\mid\mid\Phi\otimes\mathbbm{I}_{\mathcal{H}_{R}}\mid\mid_{1}$ for any $\mathcal{H}_{R}$ with $\dim(\mathcal{H})\leq\dim(\mathcal{H}_{R})$ [26]. It should be mentioned here that for a pair of arbitrary linear maps $\Phi_{1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{K}_{1})$ and $\Phi_{2}:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{K}_{2})$, the equality $\mid\mid\Phi_{1}\otimes\Phi_{2}\mid\mid_{\Diamond}=\mid\mid\Phi_{1}\mid\mid_{\Diamond}.\mid\mid\Phi_{2}\mid\mid_{\Diamond}$ holds [26]. Furthermore, if $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})$ is a quantum channel, $\mid\mid\Phi\mid\mid_{\Diamond}=1$ [26].
@@
After performing a measurement $M=\{M(x)\}$, one can post-process the outcomes which is equivalent to performing another measurement $N=N(y)$. For this purpose, if one uses the probability distributions $\nu_{x}=\{\nu_{x}(y)\}$ to post-process the outcome $x$ then
$N(y)=\sum_{x\in\Omega_{M}}\nu_{x}(y)M(x)\penalty 10000\ \forall y\in\Omega_{N}.$ (12)
A measurement $M$ can be performed using many different quantum instruments. A quantum instrument $\mathcal{I}$ is defined as a set of CP trace non-increasing linear maps $\{\Phi_{x}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{K})\}_{x\in\hat{\Omega}_{\mathcal{I}}}$ that sums up to a quantum channel $\Phi=\sum_{x}\Phi_{x}$ where $\hat{\Omega}_{\mathcal{I}}$ is the outcome set of the instrument $\mathcal{I}$ [24]. It is well-known that both quantum measurements and quantum channels are special cases of quantum instruments. If the instrument $\mathcal{I}$ implements the measurement $M\in\mathscr{M}(\mathcal{H})$, then
@@
hold for all $x\in\hat{\Omega}_{\mathcal{I}}$ and here, $\hat{\Omega}_{\mathcal{I}}=\Omega_{M}$.
@@
Clearly, the measurement $\mathcal{I}^{\dagger}[N]$ has the outcome set $\Omega_{\mathcal{I}^{\dagger}[N]}=(\hat{\Omega}_{\mathcal{I}}\times\Omega_{N})$. A general transformation of a measurement can be the compositions of postprocessing of outcomes, probabilistic mixing with another measurement and action of quantum instruments. Similar to the sets of quantum channels, transformations of a set of measurements is slightly more general. Such transformations allow the controlled implemention of measurements in that set, in general.
@@
$|\displaystyle\Gamma_{M}(\rho)=\sum_{a}Tr[\rho M(a)]\ket{a}\bra{a}$ (15)
[…]
@@
$=$ $|\displaystyle\sum_{ab}Tr[\rho_{AB}(M(a)\otimes N(b))]\ket{a}\bra{a}$
$=$ $|\displaystyle\sum_{a}Tr[\rho_{AB}(M(a)\otimes\sum_{b}N(b))]\ket{a}\bra{a}$
$=$ $|\displaystyle\sum_{a}Tr[\rho_{AB}(M(a)\otimes\mathbbm{1}_{\mathcal{H}_{B}})]\ket{a}\bra{a}$
@@
Therefore, $Tr_{\mathcal{H}_{\Omega_{N}}}$ is equivalent to the marginalization operation on the measurement $N$. Clearly, for an arbitrary measurement $M=\{M(m_{1},m_{2})\}\in\mathscr{M}(\mathcal{H})$ with outcome set $\Omega_{A}=X\times Y$ where the set $X$ and $Y$ are some sets of natural numbers and $m_{1}\in X$ and $m_{2}\in Y$ we can define a channel $\Gamma_{M}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}_{X}\otimes\mathcal{H}_{Y})$ with $\dim{\mathcal{H}_{X}}=|X|$ and $\dim{\mathcal{H}_{Y}}=|Y|$. Moreover, we denote the merginalization process as $\mathbf{Merg}_{Y}$. In other words, $\mathbf{Merg}_{Y}(M)=\{[\mathbf{Merg}_{Y}(M)](m_{1})=\sum_{m_{2}}M(m_{1},m_{2})\}$. Clearly,
$\Gamma_{\mathbf{Merg}_{Y}(M)}=(\mathbbm{I}_{\mathcal{H}_{X}}\otimes Tr_{\mathcal{H}_{Y}})\circ\Gamma_{M}.$ (17)
@@
Consider two arbitrary measurements $M=\{M(a_{1},a_{2})\}\in\mathscr{M}(\mathcal{H})$ and $N=\{N(b_{1},b_{2})\}\in\mathscr{M}(\mathcal{H})$ with $\Omega_{M}=\Omega_{N}=X\times Y$. Clearly,
@@
where $M_{CBA}=\mathbbm{I}^{\dagger}_{\mathcal{H}_{C}}\otimes\mathbf{SWAP}^{\dagger}_{\mathcal{H}_{B}\leftrightarrow\mathcal{H}_{A}}(M_{CAB})$ and $N_{CBA}=\mathbbm{I}^{\dagger}_{\mathcal{H}_{C}}\otimes\mathbf{SWAP}^{\dagger}_{\mathcal{H}_{B}\leftrightarrow\mathcal{H}_{A}}(N_{CAB})$ . In Section III.1, we will define the distance for two sets of measurements.
@@ II.2 Aspects of a resource measure based on set of quantum measurements
Certain natural assumptions are respected by any state based resource theories as defined in Ref. [27, 1]. Similarly, in case of any measurements-based resource theory, one can make the following natural assumptions-
@@
    Two objects $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are free if and only if $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is free.
@@
    Two transformations $\mathcal{W}_{1}$ and $\mathcal{W}_{2}$ are free if and only if $\mathcal{W}_{1}\otimes\mathcal{W}_{2}$ is free.
@@
It should be noted that all three known measurement-based resource theories respect the natural assumptions [A1]–[A6] and are expected to be respected by any measurement-based resource theories to be developed in the future. Merginalization is a free transformation for the resource theory of measurement incompatibility and the resource theory of measurement coherence. Now, it is known that the post-processing of outcomes is not a free transformation for the resource theory of measurement sharpness, in general [13, 14]. However, merginalization is very special case of post-processing of outcomes that transforms an arbitrary measurement $M\in\mathscr{M}(\mathcal{H})$ to $\mathcal{T}_{\mathcal{H}}$ which is a trivial object. Therefore, technically, the merginalization should be the free transformation for the the resource theory of measurement sharpness. But in Ref. [14], only the outcome set preserving free transformations for the resource theory of measurement sharpness have been considered. The assumption [A6] will not be used everywhere in the manuscript. We will explicitly mention wherever we use the assumption [A6].
@@
    (Non-negativity and faithfulness): $\mathbbm{R}(\mathcal{M})\geq 0$ for all objects $\mathcal{M}$ and $\mathbbm{R}(\mathcal{M})=0$ if and only if $\mathcal{M}\in\mathcal{F}$ .
@@
    (Monotonicity): $\mathbbm{R}(\mathcal{M})\geq\mathbbm{R}(\mathcal{W}(\mathcal{M}))$ for all objects $\mathcal{M}$ and for all free transformations $\mathcal{W}$ i.e., for all $\mathcal{W}\in\mathcal{O}$ .
@@
    (Convexity): $\mathbbm{R}(\sum^{n}_{i}p_{i}\mathcal{M}_{i})\leq\sum^{n}_{i}p_{i}\mathbbm{R}(\mathcal{M}_{i})$ for all sets of objects $\{\mathcal{M}_{i}\}^{n}_{i=1}$ and all probability distributions $\{p_{i}\}^{n}_{i=1}$ .
@@
For convex measurement-based resource theories, the resource robustness of an arbitrary set of measurements $\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
[…]
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the given set of measurements $\mathcal{M}$.
Similarly, for convex measurement-based resource theories, the resource weight of an arbitrary set of measurements $\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H}_{A})$ is defined as
[…]
Here, $\tilde{\mathcal{M}}$ is an arbitrary set of unwanted noise measurements, $\mathcal{F}_{\mathcal{H}_{A}}$ is the set of free sets of measurements acting on Hilbert space $\mathcal{H}_{A}$ and the optimization is over all variables, other than the given set of measurements $\mathcal{M}$.
@@
For any resource measure $\mathbbm{R}$ in any measurement-based resource theory, and for a generic distance $\mathbf{D}$, the corresponding $\epsilon$-measure $\mathbbm{R}^{\mathbf{D}}_{inf,\varepsilon}$ of a set of measurements $\mathcal{M}=\{M_{i}\}\subset\mathscr{M}(\mathcal{H})$ is defined as
$\mathbbm{R}^{\mathbf{D}}_{inf,\epsilon}(\mathcal{M})=\inf_{\begin{subarray}{c}\mathcal{N}\subset\mathscr{M}(\mathcal{H}),\\ \mathbf{D}(\mathcal{M},\mathcal{N})\leq\epsilon,\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N}).$ (25)